OpenStudy (anonymous):

Ready for more metals? CoCa algebra 1 help :D

OpenStudy (anonymous):

find the solution of the system y=x^2-5x+2 and y=-6x+4

OpenStudy (abb0t):

Use substitution method!

OpenStudy (anonymous):

1.(2,8) and (-1,2) 2. (-2, 8) and (1, -2) 3. (-2, 16) and (1, -2) 4. no solution

OpenStudy (abb0t):

And, you spelled \(\sf \color{red}{medals}\) incorrectly.

OpenStudy (anonymous):

Lol i know, its been a rough week

OpenStudy (abb0t):

Use substitution method for this problem. Are you familiar with that method?

OpenStudy (anonymous):

No. And i really dont have time right now :/

OpenStudy (abb0t):

Well, set them equal to each other first of all, substitute the second one: \(\sf \color{red}{-6x+4}\) to \(\sf \color{blue}{y}\), so yu will have: \(\sf \color{red}{-6x+4 = x^2-5x+2}\). Combining like terms to solve for \(\sf \color{green}{x}\)! Then once you solve for \(\sf \color{green}{x}\), you can plug it in to your either equation to solve for \(\sf \color{blue}{y}\) Try it.